OpenStudy (jessie13):

Will Medal Please answer ASAP!!! Given: x - 7 > -2. Choose the solution set. {x | x _ R, x > -9} {x | x _ R, x > -5} {x | x _ R, x > 5} {x | x _ R, x > 14}

jhonyy9 (jhonyy9):

add to both sides 7 so what will get ?

OpenStudy (jessie13):

i dont get what ur saying

jhonyy9 (jhonyy9):

x-7+7 > -2+7 added,assumed to both sides 7 ,yes ?

OpenStudy (jessie13):

ya

jhonyy9 (jhonyy9):

make the calculi please

jhonyy9 (jhonyy9):

-7+7 = ?

OpenStudy (jessie13):

0

jhonyy9 (jhonyy9):

yes sure so on the other side -2+7 = ?

OpenStudy (jessie13):

5

jhonyy9 (jhonyy9):

yes on the left side what will remain there without cancel out the +7-7 ?

OpenStudy (jessie13):

-14

jhonyy9 (jhonyy9):

no said above that -7+7 = 0 yes ? and there is again x so what will remain after you get -7+7=0 ?

OpenStudy (jessie13):

oh sorry i thought that it said -7

jhonyy9 (jhonyy9):

was there x+7-7 > -2+7 continue it please

jhonyy9 (jhonyy9):

courage please this is easy

jhonyy9 (jhonyy9):

on the left side you know that +7-7 = 0 so what will remain on the left side ?

jhonyy9 (jhonyy9):

hi Levi ! try solve it you please

OpenStudy (braydonlevi99):

Okay. So, the problem is x - 7 > -2 Correct?

jhonyy9 (jhonyy9):

yes and like a first step i ve said above that you assume to both sides 7 and so what will get ?

OpenStudy (braydonlevi99):

So, add 7 to both sides. Then, your equation becomes as follows: x > 5

jhonyy9 (jhonyy9):

exactly very nice so than you know this how you think what choice will be right answer from there ?

OpenStudy (braydonlevi99):

What does the x _ R mean?

jhonyy9 (jhonyy9):

x real number

OpenStudy (jackthegreatest):

all real numbers

OpenStudy (jackthegreatest):

basically @jessie13 you want the answer that includes all real numbers greater than 5

OpenStudy (braydonlevi99):

Isn't the answer {x | x _ R, x > 5} ?

jhonyy9 (jhonyy9):

yes Levi this is sure

OpenStudy (jackthegreatest):

yep

jhonyy9 (jhonyy9):

good luck bye bye and hope the asker understood it right from these